# QuEST: Low-bit Diffusion Model Quantization via Efficient Selective Finetuning

### v3

@@ 2.2 Quantization-aware Properties of Diffusion Models
Property : Different activations have different sensitivity towards decreasing bit-width. Low-bit activation quantization is often more difficult than weight quantization and remains the key challenge for diffusion model quantization. Compared with weights that can be categorized into linear layers and convolutional layers, activation types are more diverse and complicated. Figure 3 shows the sensitivity of different layers’ activations to quantization. Specifically, we quantize three different types of activations to lower bits while maintaining the others’ bit-width to 8-bit, and observe how the decreasing bit-width effects generation performance. Surprisingly, we observe that the FeedForward layer [8] activations cause generation failure at as large as 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits. This indicates that a small amount of activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
2.3 Quantization via Efficient Selective Finetuning
We first theoretically discuss the potential reasons for the failure of previous methods and the underlying rationale for weight fine-tuning. Then we introduce a finetuning method QuEST for diffusion models that can significantly boost low-bit performance with less time and memory usage. Figure 1(c) shows our framework, being a distillation-based finetuning strategy consisting of selective weight optimization and network-wise scaling factor optimization. To further address the loss of temporal information due to quantization, an efficient time-aware activation quantizer is proposed.
2.3.1 Sufficiency of Finetuning
Motivated by the phenomenon that current PTQ methods completely fail when applied to low-bit diffusion models and driven by the phenomenon that the performance degradation caused by activation bit-width decrease is much more significant than weight bit-width (8-bit to 4-bit) in diffusion models, we would like to explore theoretically why previous PTQ methods fail and how to overcome this.
@@ 2.3.1 Sufficiency of Finetuning
We first review the underlying theory underpinning previous methods, which typically employ the reconstruction-based post-training quantization approach. Denote the diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the task loss as $L(z_{n,t},\overline{z}_{n,t})$, where $n$ is the number of layers and $\overline{z}_{n,t}$ is the ground-truth. $L$ can be any loss function and here we use the mean squared error (MSE). We treat quantization as a type of perturbation and formulate the influence of activation quantization using Taylor expansion, assuming model weight $\mathbf{w}$ is frozen:
$\mathbb{E}[L(z_{n,t}$ $+\Delta;\mathbf{w})]-\mathbb{E}[L(z_{n,t};\mathbf{w})]\approx\Delta^{T}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (3)
@@
$\mkern-18.0mu\frac{1}{2}\Delta^{T}\overline{\mathbf{H}}^{(z_{n,t})}\Delta=\frac{1}{2}(\tilde{z}_{n,t}-z_{n,t})^{T}\overline{\mathbf{H}}^{(z_{n,t})}(\tilde{z}_{n,t}-z_{n,t}).\mkern-18.0mu$ (4)
This leads to a rather straightforward conclusion: Quantization error can be reduced by reconstructing the model output. However, under low-bit settings, the reasoning from Equation 3 to Equation 4 is inaccurate, where the activation perturbation $\Delta$ is too large for meaningful Taylor expansion. Thus we have the following proposition:
@@
To resolve the above problem, we can use the following theorem by transforming $\Delta$ into a smaller perturbation $\epsilon$:
@@
Given an $n$ layer diffusion model at time $t$ with quantized activations as $\tilde{\mathbf{z}}_{t}=[\tilde{z}_{1,t},\tilde{z}_{2,t},...,\tilde{z}_{n,t}]$ and $\tilde{z}_{n,t}=z_{n,t}+\Delta$, where $z_{n,t}$ is the ground-truth and $\Delta$ is the large perturbation caused by low-bit quantization. Denote the target task MSE loss as $L(\tilde{z}_{n,t},\overline{z}_{n,t})$, the quantization error can be transformed into:
@@
$\approx$ $2\epsilon^{T}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-z_{\text{FP}})+\frac{1}{2}\sum_{i=1}^{K}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})^{T}\overline{\mathbf{H}}^{(z_{n,t}+(i-1)\epsilon)}(\tilde{z}_{n,t}^{i}-z_{\text{FP}})$ (5)
where $\mathbf{w}_{n}$ is the weight for layer $n$ and $\tilde{z}_{n,t}^{i}=\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}$, $\epsilon$ is a perturbation small enough for Taylor expansion, $K$ is a constant and $\Delta=K\epsilon$.
Proof is provided in Appendix B. Theorem 2.2 indicates that: To minimize the quantization error, ideally we can finetune $\mathbf{w}_{n}$ so that for any $i$, the weights can fit to the data/activation $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$ which the full-precision counterpart may not fit, making the model more robust to varying inputs. In other words, we are finetuning the model weights for better robustness towards large input activation perturbations, therefore enabling easier quantization. We also find that the second term in Equation 2.2 can be minimized simultaneously with the first term, so that the total quantization error can be reduced via a common goal: Activation output alignment via finetuning $\mathbf{w}_{n}$.
2.3.2 Quantization via Selective Finetuning
Motivated by the unraveled properties of quantized diffusion models and the shortcomings in existing quantization approaches, we propose an efficient finetuning strategy for low-bit diffusion model quantization. The components of our method are detailed below.
Data-Free Partial-network-wise Training. To alleviate the need of substantial training data, we construct the calibration set in a data-free manner. By feeding sampled Gaussian noises $x_{T}$ into the full-precision model and sample iteratively over different time steps, we can obtain the calibration data needed for finetuning the quantized model. In practice, the number of samples generated for each time step is 128 or 256, indicating that we only have to inference the full-precision model a few times to obtain the needed number of calibration samples.
As depicted in Figure 1(c), to overcome the quantization difficulty demonstrated in property efficiently, we update partial model weights ($\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$) that only account for a small subset of parameters related to the time step $t$. The other weight parameters (such as $\mathbf{W}_{l^{{}^{\prime}}}$, $\mathbf{W}_{l^{{}^{\prime\prime}}}$) are frozen during optimization. For example on LDM-4 [32], at most 7% of the parameters are included for adjustment. The choices for the weights to be finetuned are discussed in the following sections. To better conserve the performance and behavior of the full-precision model, we use both its intermediate activations and final outputs as the ground-truth for supervision. Figure 4 visualizes the generation results and demonstrates the effectiveness of this strategy in preserving image details.
The activation quantization parameters can be viewed as additional model parameters, therefore we further propose a network-wise training strategy. Different from quantization methods using layer-wise or block-wise reconstruction [19, 34] that binds quantization parameters with their corresponding layers or blocks, we optimize all activation scaling factors for each time step together with the partial weight parameters. Additionally, while layer/block-wise optimization methods can only reconstruct sequentially, we update the required parameters at once. In this way, we significantly save the time and memory needed for quantization. Note that the quantization parameters for the weights are not updated once initialized, in order to maintain memory efficiency during deployment.
Efficient Time-aware Activation Quantization. As discussed in property , accurate time embedding quantization is critical. To discriminate between different time steps, we design a temporal quantizer specifically for activation quantization that adopts different sets of quantization parameters for different time steps. While these parameters consumes negligible memory and does not effect inference speed, learning independent parameters for each time step leads to immense time cost.
Motivated by the fact that neighboring time steps yield similar functions in diffusion models [3], we cluster neighboring time steps so that they share the same quantization parameters and are calculated by the same calibration data $x_{t}$: $\mathbf{s}(t)=\mathbf{s}(t+1)=...=\mathbf{s}(t+\Delta t)$, where the time steps used for optimization are sampled uniformly, and each chosen time step’s statistics are used to represent $\Delta t$ adjacent ones. Though most time steps’ quantization parameters are approximated, this strategy saves the optimization time by $t$ times without compromising much performance. Section 3.2 further provides an analysis on the performance and efficiency trade-off.
Table 3: Unconditional image generation results of LDM-4 models on LSUN-Bedrooms/Churches 256$\times$256. "N/A" denotes generation failure. "FP" denotes the full precision model.
@@ 2.3.2 Quantization via Selective Finetuning
Selective & Progressive Layer Alignment. As investigated in Section 2.2, we identify two crucial types of activations for image generation performance and select them for weight finetuning: Time embeddings and attention-related activations (e.g. FeedForward layers in Stable Diffusion). We progressively update these components due to their distinct, non-overlapping functionalities.
Figure 4: Generated images compared with Q-Diffusion under two bit-width types: W8A8 and W4A4, using LDM-4 for LSUN-Bedrooms. Our method not only succeed well in generating low-bit images, but also aligns more closely with the full-precision counterpart.
In a single forward process, identical time embeddings are injected into different parts of the model, passed through projection layers, and merged with the latent image representations. This implies that the time information operates independently from the primary network flow. Thus, we first refine the time embedding layer $l$’s weight $\mathbf{w}_{l}$ along with its activation quantization parameters $\mathbf{s}_{l}(t)$:
$\mathcal{L}_{\text{TE}}^{t}=\sum_{l\in\mathbb{C}_{\text{TE}}}\mathbb{E}_{t}[||\overline{O}(e(t);\mathbf{w}_{l})-\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))||^{2}],$ (6)
where $\mathbb{C}_{\text{TE}}$ represents the set of time embedding layers, $e(t)$ is the fixed time embedding for time $t$. $\overline{O}(e(t);\mathbf{w}_{l})$ is the final activation output of the full-precision model representing the ground-truth, and $\tilde{O}(e(t);\mathbf{w}_{l},\mathbf{s}_{l}(t))$ is the quantized output. This objective function indicates that the chosen weight parameters are consistently updated across different time steps, so as to ensure robustness to diverse temporal inputs. Nevertheless, only the small subset of scaling factors relevant to each specific time step are adjusted for temporal information discrimination.
Apart from time embedding layers, attention-related layers are also vital (according to property ). As depicted in Figure 3, reducing the activation bit-width of these layers degrades model performance the most dramatically. Denote $\mathbb{C}_{\text{A}}$ as the set containing all these layers and given the image calibration samples $x_{t}$, we optimize the corresponding weights $\mathbf{w}_{l}$ and quantization parameters $\mathbf{s}(t)$. However, we exclude the quantization parameters already optimized in Equation 6:
$\mkern-18.0mu\mathcal{L}_{\text{A}}^{t}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||\overline{O}(x_{t};\mathbf{w}_{:l})-\tilde{O}(x_{t};\mathbf{w}_{:l},\mathbf{s}(t)\setminus s_{\mathbb{C}_{\text{TE}}}^{t})||^{2}],\mkern-18.0mu$ (7)
where $\mathbf{w}_{:l}$ are the weight parameters of the first $l$ layers. This strategy resembles linear probing, where only the layer directly associated with the activation output is finetuned.
Furthermore, we aim not only for each independent layer’s quantization error reduction, but also for the final generated images’ quality improvement. Therefore, we adopt the Task Loss Guidance strategy where the target task loss $\sum_{t}\mathcal{L}_{d}^{t}$ is introduced and the ground-truth is approximated by the output of the full-precision counterpart. By integrating Equation 6 and Equation 7, the final objective is formulated as follows:
$\mathcal{L}$ $=\sum_{t}((\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{d}^{t})+(\mathcal{L}_{\text{A}}^{t}+\mathcal{L}_{d}^{t}))$
$=\sum_{t}(\mathcal{L}_{\text{TE}}^{t}+\mathcal{L}_{\text{A}}^{t}+2\mathbb{E}[||\overline{O}(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s}(t))||^{2}]).$ (8)
@@ 3.1 Experiment Settings
To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets [39], class-conditional image generation on ImageNet [5], and text-to-image generation. The model architectures we quantize include LDM and Stable Diffusion [32], and uses "WnAm" to represent the quantization setting: n-bit weight quantization and m-bit activation quantization. DDIM samplers [15] are adopted for LDMs and the PLMS sampler [26] is used for Stable Diffusion. We generate 256 samples per time step for constructing the calibration set, and train 20 epochs over each time step. The Adam optimizer [18] is adopted and the learning rate for weight finetuning and scaling factor finetuning are set as $1e^{-5}$ and $1e^{-4}$ respectively.
Table 4: Class-conditional generation results using LDM-4 on ImageNet 256$\times$256.
@@ 3.2 Experiment Results and Analysis
Unconditional Generation: We evaluate the performance of our method over LDM-4 (LSUN-Bedrooms 256$\times$256) and LDM-8 (LSUN-Churches 256$\times$256) using the DDIM sampler with 200 and 500 time steps, respectively. Results are shown in table 3. The results are evaluated using FID, where the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet [19].
Table 5: Text-to-image generation on Stable Diffusion (512$\times$512) using COCO2014 prompts.
@@
Class-conditional Generation: We evaluate the performance using LDM-4 on ImageNet 256$\times$256 using the DDIM sampler (20 steps). As shown in Table 4, three metrics are used for evaluation. Note that sFID uses additional intermediate spatial features for calculation compared with FID. We can also see that FID is not a valid metric for ImageNet LDM-4 evaluation: All methods have lower FID when quantized to lower-bits, conflicting with human perception. We show that our method not only succeeds in W4A4 quantization, but also significantly improve the generation quality on higher bit settings. Under W8A8, our method is able to achieve 2.82 improvement over sFID and 8.13 increase in Inception Score. Under W4A8, we improve the SOTA PTQ method by 1.57 sFID and 8.77 IS.
Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps. Table 5 shows the results. Images are generated based on the 10,000 prompts sampled from the COCO2014 [24] validation set, and CLIP Score is calculated based on the ViT-B/16 backbone. Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
Table 6: Influence of task loss supervision on model performance.
@@
Efficiency comparison and effects of components. Table 6 compares with and without the use of task loss $\mathcal{L}_{d}$. The results show that using the difference between the outputs of the quantized model and the full-precision counterpart for supervision is essential for performance improvement, enhancing the FID by 2.58 and 5.21 for $\text{SLA}_{\text{TE}}$ and $\text{SLA}_{\text{A}}$, respectively. However, when the learning process is only supervised by the task loss, the performance degrades by 7.13 FID and 9.39 sFID for $\text{SLA}_{\text{TE}}$, indicating that the task loss alone is insufficient.

### v4

@@ 3.3.3 Critical Module Alignment
Fig. 3 illustrates the sensitivity of different activations to quantization. Specifically, we quantize three different types of activations to lower bits while maintaining the others’ bit-width to 8-bit, and observe how the decreasing bit-width affects generation performance. Compared to weights that only fall into linear and convolutional layers, activations are more diverse and complex, making their effective quantization more challenging. Surprisingly, we observe that the FeedForward layer [7] activations cause generation failure at as early as 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits. This indicates that these activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
Denote $\mathbb{C}_{\text{A}}$ as the set containing all attention-related layers and given their image calibration inputs $z_{t,l}$, we optimize the corresponding weights and all quantization parameters $\mathbf{s}$, except for the ones already updated:
$\mathcal{L}_{\text{CMA}}=\sum_{l\in\mathbb{C}_{A}}\mathbb{E}_{t}[||O(z_{t,l};\mathbf{w}_{l})-\tilde{O}(\tilde{z}_{t,l};\mathbf{w}_{l},\mathbf{\hat{s}})||^{2}],$ (5)
where $\mathbf{w}_{l}$ are the weight parameters of the $l$th layer, $\tilde{z}_{t,l}$ is the quantized layer input, $\mathbf{\hat{s}}=\mathbf{s}\setminus\mathbf{s}_{l},l\in\mathbb{C}_{\text{TE}}$, which represents all the quantization parameters without the ones already finetuned in Eq. 4. Note that we use different inputs to optimize each module, so as to enhance the robustness of the modules to the input perturbations.
3.3.4 Progressive Alignment with Global Loss
As investigated in the previous sections, two crucial types of layers are identified and selected for weight finetuning to enable quantization efficiency: time embedding layers and attention-related layers. We progressively align these components with the full-precision model due to their distinct, non-overlapping functionalities. Since temporal information is independent of the image input and determined early in the model, we first finetune the time embedding layers to provide accurate time step guidance for each subsequent module. Then we optimize the attention-related modules with the refined time embeddings.
However, the above selective finetuning strategy only aligns the local information in the model, but is unaware of the global error reduction of the quantized model and the quantization parameters of the unselected layers. To improve the final generated images’ quality, we further aim to minimize the target task loss to provide global supervision:
$\mathcal{L}_{\text{G}}=\mathbb{E}_{t}[||O(x_{t};\mathbf{w})-\tilde{O}(x_{t};\mathbf{w},\mathbf{s})||^{2}],$ (6)
where $\mathbf{w}$ represents all the model weights, $O(x_{t};\mathbf{w})$ represents the final output of the full-precision counterpart and $\tilde{O}(x_{t};\mathbf{w},\mathbf{s})$ is the final output of the quantized model.
By integrating Eq. 4, Eq. 5 and Eq. 6, the final objective is formulated as:
$\mkern 28.0mu\arg\min_{\mathbf{w}_{l}}(\mathcal{L}_{\text{TLA}}+\mathcal{L}_{\text{CMA}}+2\mathcal{L}_{\text{G}}),\ \ l\in\mathbb{C}_{\text{TE}}\cup\mathbb{C}_{\text{A}}.$ (7)
3.4 Finetuning from a Theoretical Perspective
The above proposed method is motivated by the intuition that finetuning the model weights can adjust the activation distribution such that the imbalanced activation phenomenon can be alleviated. In this part, we attempt to explain why finetuning may be a feasible solution, offering additional insights for readers. However, we note that this is not a theoretical guarantee of the proposed method.
@@ 3.4 Finetuning from a Theoretical Perspective
We first review the underlying theory underpinning conventional post-training-quantization methods, which typically employ the reconstruction-based approach. Denote the full-precision diffusion model’s activations at time $t$ as $\mathbf{z}_{t}=[z_{1,t},z_{2,t},...,z_{n,t}]$, the final loss as $L(\mathbf{z}_{t};\mathbf{w})$, where $n$ is the number of layers. $L$ can be any loss function and here we use the mean squared error (MSE). We treat quantization as a type of perturbation and formulate the influence of activation quantization using Taylor expansion, assuming model weight $\mathbf{w}$ is frozen:
$\approx\Delta^{\mathrm{T}}\overline{\mathbf{g}}^{(z_{n,t})}+\frac{1}{2}\Delta^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t})}\Delta,$ (8)
@@
$\mkern-18.0mu\frac{1}{2}\Delta^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t})}\Delta=\frac{1}{2}(\tilde{z}_{n,t}-z_{n,t})^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t})}(\tilde{z}_{n,t}-z_{n,t}).\mkern-18.0mu$ (9)
However, under low-bit settings, the reasoning from Sec. 3.4 to Eq. 9 is inaccurate, where the activation perturbation $\Delta$ is too large for a meaningful Taylor expansion. Thus we have the following proposition:
@@
Since the inaccuracy arises from the large activation perturbation $\Delta$, we transform $\Delta$ into a smaller perturbation $\epsilon$ and derive the following theorem:
@@
Given an $n$ layer diffusion model at time $t$ with quantized activations as $\tilde{\mathbf{z}}_{t}=[\tilde{z}_{1,t},\tilde{z}_{2,t},...,\tilde{z}_{n,t}]$ and $\tilde{z}_{n,t}=z_{n,t}+\Delta$, where $z_{n,t}$ is the ground truth and $\Delta$ is the large perturbation caused by low-bit quantization. Denote the target task MSE loss as $L(\mathbf{z}_{t};\mathbf{w})$, the quantization error can be transformed into:
@@
$\approx$ $2\epsilon^{\mathrm{T}}\sum_{i=1}^{K}(\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}-z_{\text{n,t}})$
$+\frac{1}{2}\sum_{i=1}^{K}(\tilde{z}_{n,t}^{i}-z_{\text{n,t}})^{\mathrm{T}}\overline{\mathbf{H}}^{(z_{n,t}+(i-1)\epsilon)}(\tilde{z}_{n,t}^{i}-z_{\text{n,t}})$ (10)
where $\mathbf{w}_{n}$ is the weight for layer $n$ and $\tilde{z}_{n,t}^{i}=\tilde{z}_{n-1,t}^{i}\cdot\mathbf{w}_{n}$, $K$ is a large constant and $\Delta=K\epsilon$.
@@
Table 3: Quantization performance on LSUN-Bedrooms/Churches 256$\times$256. “N/A” denotes generation failure. “*” denotes the results obtained by re-implementing the open-source code. More baseline and metric comparisons are included in the Appendix.
Theorem 3.2 indicates that, to minimize quantization error, $\mathbf{w}_{n}$ should ideally be fine-tuned so that, for any $i$, the weights fit the corresponding input $\tilde{z}_{n-1,t}^{i}+(i-1)\epsilon$. This adjustment captures variations that the full-precision model may overlook. In other words, fine-tuning optimizes model weights for better robustness towards large input activation perturbations, facilitating easier quantization. Moreover, since the finetuned and quantized model is aligned with the original full-precision model, the potential impact on generation performance can be avoided. Note that the second term in Theorem 3.2 can be ignored within an acceptable upper bound, as it is of second order and shares a common zero-loss solution with the first term.
@@ 4.1 Experiment Settings
To verify the effectiveness of our proposed method, we conduct experiments on three types of generation tasks: Unconditional image generation on LSUN-Bedrooms and LSUN-Churches datasets [37], class-conditional image generation on ImageNet [4], and text-to-image generation. The model architectures we quantize include LDMs and Stable Diffusion [28], and use ”WnAm” to represent the quantization setting: n-bit weight quantization and m-bit activation quantization. DDIM samplers [13] are adopted for LDMs and the PLMS sampler [25] is used for Stable Diffusion. We generate 256 samples per time step for constructing the calibration set. The Adam optimizer [17] is adopted and the learning rate for weight finetuning and scaling factor finetuning is set as $1e^{-5}$ and $1e^{-4}$ respectively.
@@
Table 4: Quantization performance on ImageNet 256$\times$256. “*” denotes the results obtained by re-running the open-source code.
@@ 4.2 Experiment Results and Analysis
Unconditional Generation: We evaluate the performance of our method over LDM-4 (LSUN-Bedrooms 256$\times$256) and LDM-8 (LSUN-Churches 256$\times$256) using the DDIM sampler with 200 and 500 time steps, respectively. Results are shown in Tab. 3 using FID, where our method outperforms the other baselines by a good margin. Note that the Inception Score is not a reasonable metric for datasets that have significantly different domains and categories from ImageNet [18], thus not included. We further provide comparison with TFMQ-DM [15] in Appendix 6.
@@
Table 5: Quantization performance on Stable Diffusion v1.4 (512$\times$512) using COCO2014 prompts.
@@
Class-conditional Generation: We evaluate the performance using LDM-4 on ImageNet 256$\times$256 using the DDIM sampler (20 steps). As shown in Tab. 4, three metrics are used for evaluation. Note that sFID uses additional intermediate spatial features for calculation compared with FID. We can also see that FID is not a valid metric for ImageNet LDM-4 evaluation: All methods have lower FID when quantized to lower bits, conflicting with human perception. We show that our method not only succeeds in W4A4 quantization, but also improves the generation quality under higher bit settings. Under all three kinds of bit-width settings, our method is able to outperform the SOTA PTQ methods and EfficientDM in both sFID and IS. Examples of our generated images are included in Appendix 11.
Text-to-image Generation: We use Stable Diffusion v1.4 as the model for quantization with the PLMS sampler sampling 50 time steps. Tab. 5 shows the results. Images are generated based on the 10,000 prompts sampled from the COCO2014 [23] validation set, and CLIP Score is calculated based on the ViT-B/16 backbone. Given the limited works done on Stable Diffusion, we can only compare with Q-Diffusion and the full-precision baseline.
Figure 4: Visual comparison with Q-Diffusion and EfficientDM. QuEST outperforms the baselines with better visual quality.
4.3 Ablations and Discussions